OpenStudy (anonymous):

Find the point in the first quadrant that lies on the hyperbola y^2-x^2=5 and is closest to point (1,0). If we write the point as (a,b), then a=___ & b=___

OpenStudy (anonymous):

a is not 0-3

OpenStudy (anonymous):

if you find a, i can get b

OpenStudy (amistre64):

the slope at any point is defined as x/y i can tell that much so far

OpenStudy (amistre64):

which means the slope between our given point and the nearest point on the hyper is -y/x

OpenStudy (amistre64):

if im thinking this right, then x=1/2

OpenStudy (amistre64):

\[(\frac{1}{2},\frac{\sqrt{19}}{2})\] if my idea is correct

OpenStudy (anonymous):

we could also try this maybe would work. solve for y and get \[y^2=x^2+5\] \[y=\sqrt{x^2+5}\] and the square of the distance between \[(x, \sqrt{x^2+5}\] and \[(1,0)\] is \[(x-1)^2+\sqrt{x^2+5}^2=(x-1)^2+x^2+5\]

OpenStudy (amistre64):

y^2-x^2=5 2y y' -2x = 0 y' = x/y slope from 1,0 to x,y 1 ,0 -x-y ----- 1-x,-y -y -y ---- = ---- 1-x x 1-x=x 1=2x 1/2 = x

OpenStudy (anonymous):

\[D^2=2x^2-2x+6\] minimum of quadratic at \[-\frac{b}{2a}=\frac{1}{2}\] so i guess we get the same answer and as usual more than one way to skin a cat

OpenStudy (anonymous):

im a little confused by how you got the 1/2 but thanks